OpenStudy (anonymous):

limθ tends to 0[sinθ-θ]/[θ-tanθ]

OpenStudy (anonymous):

it Is -1

myininaya (myininaya):

\[\frac{\sin (\theta)-\theta}{\theta-\tan(\theta)}*\frac{\frac{1}{\theta}}{\frac{1}{\theta}}=\frac{\frac{\sin(\theta)}{\theta}-1}{1-\frac{\tan(\theta)}{\theta}}\] \[=\frac{\frac{\sin(\theta)}{\theta}-1}{1-\frac{\sin(\theta)}{\cos(\theta)*\theta}}\] so we 0/0 we should had used l'hospital way

myininaya (myininaya):

\[\frac{\cos (\theta)}{-\sec^2(\theta)}=-\cos^3(\theta)\]

myininaya (myininaya):

now plug in 0 in you will get -1

OpenStudy (anonymous):

it is olready in 0/0 form ... jus use lhopitals rule once nd substitute Theta =0

myininaya (myininaya):

i know i seen that after i got that far lol

OpenStudy (anonymous):

ya happens some time :P LoL

myininaya (myininaya):

yep yep

OpenStudy (anonymous):

ok.